Grade 11General Physics

two particles moving along the same straight line starting at the same moment from the same point in the same direction. The first moves with a constant velocity u and the socind with constant acceleration f. During the time that elapses before second catches the first, the greatest distance between the particles is?.....{hint: the answer is (u.u)/2f

Radhika Batra
12 Years agoGrade 11

1 Answer

shubham
8 Years ago
let x distance covered by time t
1 particle x1=vt
2 particle x2=ft^2/2
x=x1-x2
=vt-ft^2/2
for x max.
dx/dt=0
=v-ft=0
t=v/f
x=v(v/f)-fv^2/2f^2
=v^2/2f